# From Models to Network Topologies: A Topology Inference Attack in Decentralized Federated Learning

### v1

@@ 6 Evaluation
This work empirically validates the effectiveness of the proposed topology inference attack. First, it identifies which metrics yield stronger attack performance. Building on these results, it then thoroughly evaluates the attack strategy under various conditions, including different datasets, topologies, node counts, and data preprocessing methods. This comprehensive assessment demonstrates the robustness and applicability of the proposed topology inference attack.
@@ 6.1 Experimental Setups
A. Dataset and Model: All experiments are conducted on three datasets: MNIST, Fashion-MNIST, and CIFAR10, widely used in ML and inference attack benchmarking.
  • MNIST LeCun and Cortes (2010): Contains 60,000 grayscale training images and 10,000 test images of handwritten digits (0–9) at 28$\times$28 resolution. A two-hidden-layer MLP (256 and 128 neurons) trained with Adam (lr=1e-3) is used.
  • FMNIST Xiao et al. (2017): Similar size and format to MNIST, but with 10 categories of clothing items. A CNN with two convolutional layers (32 and 64 filters, kernel size 3$\times$3), ReLU, and max-pooling is employed.
  • CIFAR10 Krizhevsky et al. (2009): Contains 60,000 32$\times$32 RGB images from 10 categories. The training set is only treated with basic image normalization. The image pixel values are normalized to keep the original image content unchanged. A MobileNet Sandler et al. (2019) is used for the model training.
B. Topology Setting: All experiments employ a DFL network that includes five topologies: a ring, a star, and three Erdős-Rényi (ER) graphs with probabilities $p=0.3$, $p=0.5$, and $p=0.7$ Erdős et al. (2012). These configurations cover a spectrum of network densities, ranging from sparse to dense.
@@ C. Federation Setting: All experiments employ the following DFL configurations:
  • Number of Nodes: The network size is set to 10, 20, or 30 nodes, with training data randomly assigned to each node in an Independent and Identically Distributed (IID) manner, ensuring no overlap between nodes.
  • Total Rounds: The total number of rounds is 10, 30, and 40 for networks with 10, 20, and 30 nodes, respectively.
@@
D. Attack Models: The topology inference attack models are categorized by the attack strategy. For the supervised attack strategy, the models used are Logistic Regression Wright (1995), Support Vector Machine (SVM) Hearst et al. (1998), and Random Forest (RF) Breiman (2001). For the unsupervised attack strategy, the models include K-Means Lloyd (1982), Gaussian Mixture Model (GMM) Reynolds and others (2009), and Spectral Clustering Von Luxburg (2007).
@@ 6.2 Selection of Attack Metrics
Figure 1: Normalized distributions of attack metrics in a 10-node ring-topology DFL system trained on the MNIST dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection.
Selecting appropriate attack metrics is critical to the success of inference attacks. An effective attack metric should distinguish between connected and non-connected nodes as clearly as possible. To this end, the first part of this experiment examines the effectiveness of the four proposed attack metrics by analyzing their respective distributions. Figure 1 illustrates the distributions of these metrics for a 10-node ring topology DFL system trained on the MNIST dataset, comparing directly connected nodes (the edge group) with those not (the non-edge group). To facilitate direct comparison, all values have been normalized. The results indicate that the relative loss for connected nodes is notably lower than that for non-connected nodes, with minimal overlap between the respective distributions. By contrast, the relative entropy metric exhibits substantially greater overlap. Since computing the relative loss requires access to local models and data, it has been employed in attack scenarios 1 and 3.
Regarding similarity-based metrics, the Euclidean similarity between connected models is distinctly higher than that between non-connected models, displaying no overlap between the distributions. The cosine similarity, on the other hand, shows a considerably more significant degree of overlap. Thus, the Euclidean similarity has been adopted as the attack metric in scenarios where only local models are available (attack scenarios 2 and 4).
6.3 Selection of Attack Models
Building on the attacker’s knowledge and capabilities, the topology inference attack is formulated either as a supervised learning problem (attack scenarios 1 and 2) or an unsupervised learning problem (attack scenarios 3 and 4). Corresponding attack algorithms are developed for each scenario. Using the established attack metrics, experiments are conducted on three datasets under a DFL system with an ER 0.5 topology, aiming to select the effective attack model.
Figure 2: F1-Scores of three supervised attack models for topology inference attacks, using relative loss and Euclidean similarity on three datasets under a 10 nodes DFL system with ER 0.5 topology.
Figure 2 compares the F1-Score of three supervised algorithms across the three datasets. The results demonstrate that the Logit algorithm consistently achieves strong and stable performance in both attack scenarios and on all three datasets. Therefore, the Logit algorithm is chosen for supervised attacks (attack scenarios 1 and 2).
Figure 3 presents the F1-Score of three unsupervised algorithms evaluated on the three datasets for attack scenarios 3 and 4. Among these methods, the K-Means algorithm delivers the best and most stable results, making it the algorithm of choice for unsupervised attacks (attack scenarios 3 and 4).
Figure 3: F1-Scores of three unsupervised attack models for topology inference attacks, using relative loss and Euclidean similarity on three datasets under a 10 nodes DFL system with ER 0.5 topology.
@@ 6.4 Attack Performance
After establishing the attack metrics and selecting the appropriate attack models, the proposed topology inference attacks were evaluated on three datasets, encompassing four attack scenarios and five types of topologies. The results in Table 1 indicate that the proposed methodology is highly effective. In a 10-node network, all four attack scenarios achieve an F1-Score exceeding 0.8, demonstrating that even with access restricted to public models, there is a high probability of successfully inferring the sensitive topology of a DFL system.
@@
Table 1: Topology inference attack performance on three datasets, encompassing four attack scenarios and five types of topologies.
NODES DATASET SC1 SC2 SC3 SC4
10 CIFAR10 0.922 $\pm$ 0.062 0.874 $\pm$ 0.033 0.989 $\pm$ 0.016 0.832 $\pm$ 0.099
FMNIST 0.892 $\pm$ 0.109 0.885 $\pm$ 0.084 0.980 $\pm$ 0.044 0.849 $\pm$ 0.124
MNIST 0.852 $\pm$ 0.096 0.899 $\pm$ 0.098 0.869 $\pm$ 0.087 0.849 $\pm$ 0.124
20 CIFAR10 0.877 $\pm$ 0.080 0.749 $\pm$ 0.068 0.843 $\pm$ 0.115 0.584 $\pm$ 0.166
FMNIST 0.822 $\pm$ 0.115 0.769 $\pm$ 0.103 0.748 $\pm$ 0.151 0.607 $\pm$ 0.153
MNIST 0.759 $\pm$ 0.126 0.712 $\pm$ 0.260 0.651 $\pm$ 0.178 0.698 $\pm$ 0.147
30 CIFAR10 0.815 $\pm$ 0.144 0.719 $\pm$ 0.155 0.800 $\pm$ 0.154 0.534 $\pm$ 0.213
FMNIST 0.731 $\pm$ 0.156 0.787 $\pm$ 0.097 0.697 $\pm$ 0.183 0.604 $\pm$ 0.205
MNIST 0.683 $\pm$ 0.148 0.809 $\pm$ 0.114 0.439 $\pm$ 0.134 0.615 $\pm$ 0.182
Generally, attack scenario 1 yields the highest attack success rate when the number of nodes equals 20 and 30 due to its access to the most information, including local models, data, and partial edge knowledge. Meanwhile, the performance in attack scenario 2 declines when local data are unavailable, underscoring the critical importance of local data for accurate topology inference. Notably, attack scenario 3 also achieved a high attack success rate despite lacking edge information. This finding suggests that, compared to edge knowledge, access to local data contributes more to the successful inference of the topology in DFL. The complete set of experimental results is available in the supplementary materials.
A. Effect of Topology Density. To investigate the influence of topological characteristics on topology inference attacks, Figure 4 presents the F1-Scores achieved by attack scenario 1 and attack scenario 3 under five different topologies with 20 nodes(complete experimental results are provided in the supplementary materials). Among these, the star topology consistently achieves a high inference success rate. This is because each leaf node is connected solely to the central node in the star topology, thereby amplifying the differences in model parameters between leaf nodes and making the topology particularly easy to infer.
Additionally, under attack scenario 1, the inference performance negatively correlates with network density. In denser networks (such as ER 0.7), frequent information exchange between nodes reduces model heterogeneity, limiting the supervised attack model’s generalization and inference effectiveness. In contrast, network density does not adversely affect the unsupervised model.
Figure 4: F1-Scores achieved by attack scenario 1 and attack scenario 3 under five different topologies with 10 nodes.
B. Effect of Network Size. To evaluate the scalability of the proposed topology inference attacks, experiments were extended to networks comprising 20 and 30 nodes. The results, summarized in Table 1, indicate that increasing the number of nodes generally makes the inference task more challenging, leading to a decline in attack performance. Nevertheless, under attack scenarios 1 and 2, F1-Scores consistently remain above 0.8 across all three datasets, demonstrating robust scalability. In contrast, attack scenario 4 exhibits a more pronounced decrease in performance as the network size increases.
@@ 7.1 Reduce Overfitting
Similar to other inference attacks, the success of topology inference attacks also relies on the overfitting of the target model. To validate this, an experiment was conducted on a 10-node CIFAR10 network by increasing the number of local training epochs to induce greater overfitting. As shown in Figure 5, raising the local training epochs from 3 to 10 enhances the attack success rate under both attack scenario 1 and attack scenario 3 across all examined topologies. This result indicates that overfitting intensifies the risk of information leakage in DFL systems and suggests that mitigating overfitting can effectively mitigate topology inference attacks.
In line with this, data augmentation techniques were employed in the CIFAR10 dataset to mitigate overfitting. Specifically, each training image was randomly cropped by 4 pixels and padded. As depicted in Figure 5, introducing data augmentation substantially decreases the F1-Scores of the attacks across various topologies. Notably, for attack scenario 3 in the ER 0.5 network, the F1-score drops by as much as 50%. These results demonstrate that data augmentation not only enhances model generalization but also mitigates the risk of information leakage.
@@ 7.2 Data Heterogeneity
The distribution of data across nodes significantly influences the DFL model performance. Previous experiments have been conducted under the IID condition, where data is uniformly allocated across nodes. To explore the effects of data heterogeneity, specifically non-IID distributions, a Dirichlet distribution with a parameter of $\alpha$ = 0.1 was utilized to simulate these non-IID conditions. Experiments were carried out using a CIFAR10 dataset distributed across 10 nodes, with the findings illustrated in Figure 5. The results indicate that an increase in data heterogeneity, achieved by setting the non-IID parameter $\alpha$ = 0.1, significantly diminishes the effectiveness of topology inference attacks across all tested attack scenarios and network topologies. This outcome implies that enhancing data heterogeneity may serve as an effective strategy for reducing the risk of sensitive information exposure in DFL systems.
@@ 8 Conclusion and Future Work
This work introduces a novel topology inference attack against DFL, exposing critical vulnerabilities related to privacy and information leakage. By analyzing local models, attackers can accurately infer the overlay topology, one of DFL’s most sensitive assets, highlighting the system’s susceptibility to privacy breaches. The study explores various attack scenarios and develops tailored metrics, models, and algorithms, with experiments confirming the feasibility and effectiveness of these attacks.
Furthermore, factors like network size, density, model overfitting, and data heterogeneity significantly influence attack success. Mitigation strategies, such as data augmentation and increasing data heterogeneity, can enhance model generalization and reduce the risk of information leakage.
The present work focuses on attack scenarios 1 through 4 and does not yet address the most challenging attack scenario 5. Future research plans include exploring feasible strategies for Scenario 5, as well as extending the evaluations to a wider range of datasets.

### v2

@@ 6 Evaluation
This work empirically validates the effectiveness of the proposed topology inference attack. First, it identifies which metrics yield stronger attack performance. Building on these results, it then thoroughly evaluates the proposed attack strategies under various conditions, including different datasets, topologies, node counts, and data preprocessing methods. This comprehensive assessment demonstrates the robustness and applicability of the proposed topology inference attack.
@@ 6.1 Experimental Setups
A. Dataset and Model: All experiments are conducted on five datasets: MNIST, Fashion-MNIST, CIFAR10, SVHM, and ImageNet100, widely used in FL and inference attack benchmarks.
  • MNIST [23]: Contains 60,000 grayscale training images and 10,000 test images of handwritten digits (0–9) at 28$\times$28 resolution. A two-hidden-layer MLP (256 and 128 neurons) trained with Adam (lr=1e-3) is used.
  • FMNIST [35]: Similar size and format to MNIST, but with 10 categories of clothing items. A CNN with two convolutional layers (32 and 64 filters, kernel size 3$\times$3) is employed.
  • CIFAR10 [22]: Contains 60,000 32$\times$32 RGB images from 10 categories. The training set is only treated with basic image normalization. The image pixel values are normalized to keep the original image content unchanged. A MobileNet [30] is used for training.
  • SVHN [26]: Composes of over 600,000 32$\times$32 RGB images of house numbers collected from Google Street View. Each image contains a centered digit from 0 to 9. A ResNet9 [17] is adopted as the training model.
  • ImageNet10: A curated subset of the ImageNet [8], containing 10 classes selected from the original ImageNet hierarchy. This version [1] includes approximately 15,000 images with varied object categories and balanced class distributions. A lightweight PoolFormer-S12 [37] model is used for training.
B. DFL Topology Setting: All experiments are conducted on DFL systems comprising two categories of topologies: synthetic and real-world. The synthetic topologies include ring, star, and three Erdős-Rényi (ER) graphs with connection probabilities $p=$0.3, 0.5, and 0.7 [9]. The real-world topologies consist of 29 network topologies collected from the SNDlib [27] and DEFO [16] repositories, which reflect realistic communication and infrastructure networks. Detailed descriptions and properties of these topologies are provided in the supplementary material.
@@ C. Federation Setting: All experiments employ the following DFL configurations:
  • Number of Nodes: For synthetic topologies, the number of nodes is set to 10, 20, and 30. For real-world topologies, the number of nodes ranges from 12 to 161, depending on the specific topology.
  • Total Rounds: The total number of communication rounds is determined by the rule rounds = number of nodes. The mechanism ensures efficient inter-node communication, thereby enabling consistent scalability of training time as the network size increases.
@@
  • Data Distribution: Both Independent and Identically Distributed (IID) and non-IID $\alpha=0.1$ data distributions are considered to evaluate the impact of data heterogeneity.
D. Attack Models: To evaluate the effectiveness of the proposed methods, EDGEPRE is compared against classical supervised classifiers, including Logistic Regression [34], Support Vector Machine (SVM) [18], and Random Forest (RF) [4]. For the unsupervised attack strategy, INFERGAT is evaluated against standard clustering-based methods, including K-Means [24], Gaussian Mixture Model (GMM) [28], and Spectral Clustering [32].
@@ 6.2 Selection of Attack Metrics
Figure 1: Normalized distributions of attack metrics in a 20-node star-topology DFL system trained on the FMNIST dataset. The “edge group” corresponds to metrics calculated between directly connected nodes, while the “non-edge group” corresponds to metrics calculated between nodes with no direct connection.
Selecting appropriate attack metrics is critical to the success of inference attacks. An effective attack metric should distinguish between connected and non-connected nodes as clearly as possible. To this end, the first part of this experiment examines the effectiveness of the six proposed attack metrics by analyzing their respective distributions.
Figure 1 illustrates the distributions of these metrics for a 20-node star topology DFL system trained on the FMNIST dataset, comparing directly connected nodes (edge group) with those not (non-edge group). To facilitate direct comparison, values have been normalized.
The results indicate that the relative loss for connected nodes is notably lower than that for non-connected nodes, with minimal overlap between the respective distributions. By contrast, the relative entropy and sensitivity metrics exhibit substantially greater overlap. Since computing the relative loss requires access to local models and data, it has been employed in attack scenarios 1 and 3.
Regarding similarity-based metrics, the cosine similarity between connected models is distinctly higher than that between non-connected models, displaying no overlap between the distributions. On the other hand, the curvature divergence and Euclidean similarity show a considerably more significant degree of overlap. Thus, the cosine similarity has been adopted as the attack metric in scenarios where only local models are available (attack scenarios 2 and 4). The complete set of experimental results for all datasets and topologies is in the Appendix.
@@ 6.3 Attack Performance
After establishing the attack metrics, the proposed topology inference attacks were evaluated on five datasets, encompassing four attack scenarios and different topologies. The results in Table 1 demonstrate that the proposed methods, EDGEPRE and INFERGAT, are highly effective, achieving the best performance across all four attack scenarios and five datasets.
@@
Table 1: Averaged Topology inference performance (F1-Score) under four attack scenarios across five datasets.
Scenario Algorithm CIFAR10 FMNIST ImageNet10 MNIST SVHN
SC1 Logistic 0.704 0.773 0.752 0.795 0.788
RF 0.844 0.845 0.879 0.844 0.851
SVM 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.851 0.875 0.840 0.876 0.874
SC2 Logistic 0.704 0.771 0.752 0.794 0.787
RF 0.844 0.845 0.879 0.844 0.851
SVC 0.718 0.751 0.744 0.772 0.765
EDGEPRE 0.850 0.875 0.843 0.876 0.872
SC3 GMM 0.723 0.443 0.482 0.669 0.679
Kmeans 0.727 0.715 0.618 0.724 0.734
Spectral 0.597 0.638 0.591 0.607 0.648
INFERGAT 0.800 0.788 0.733 0.799 0.802
SC4 GMM 0.576 0.662 0.685 0.670 0.688
Kmeans 0.730 0.776 0.741 0.785 0.795
Spectral 0.683 0.645 0.636 0.637 0.627
INFERGAT 0.797 0.810 0.743 0.807 0.812
It can also be observed that from SC1 to SC4, the attacker’s accessible knowledge gradually decreases, leading to a corresponding decline in attack performance. In SC1, the inferred topologies achieve an F-score exceeding 85% on all five datasets, whereas in SC4, the score drops to approximately 80%. This indicates that prior knowledge has a significant impact on the success rate of topology inference. Comparing SC1 and SC2, the difference in performance is relatively small, suggesting that combining data and model knowledge offers limited additional benefit over model-only knowledge. This implies that the model itself already encodes sufficient information to support effective topology inference.
A. Effect of Topology Density. To investigate the influence of topological characteristics on topology inference attacks, Figure 2 presents the F1-Scores achieved by each attack scenario under five in the CIFAR10 dataset. Topology density represents the ratio of the actual number of edges to the maximum possible number of edges, calculated as $\frac{2E}{N(N-1)}$. This metric reflects the overall connectivity or sparsity of a graph.
As shown in Figure 2, across four different attack scenarios, a consistent trend was observed: as the graph’s density increases, the attack success rate of the inference models (EDGEPRE and INFERGAT) tends to decrease. This is primarily because denser topologies lead to increased information fusion, resulting in higher similarity among local models. In highly dense networks, the models become increasingly homogeneous, making it more difficult for the inference algorithms to distinguish whether a connection exists between nodes. In contrast, sparse topologies exhibit greater heterogeneity among nodes despite being limited by data sparsity and class imbalance. This heterogeneity simplifies the inference task, allowing the attack models to make more accurate predictions.
Figure 2: Impact of topology density on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
B. Effect of Network Size. To evaluate the scalability of the proposed inference algorithms under different attack scenarios, Figure 3 illustrates the performance variation across four attack scenarios as the network size increases. The results show a consistent upward trend in attack effectiveness with increasing numbers of nodes, demonstrating the strong scalability of the proposed methods. On one hand, the algorithms accurately infer the underlying topology of DFL networks across varying levels of density, indicating their robustness in handling data imbalance challenges. On the other hand, even in large-scale networks with up to 161 nodes, the inference models maintain strong and stable performance, further validating the stability and scalability of the proposed topology inference approaches.
Figure 3: Impact of network size on the F1-scores achieved by each attack scenario on the CIFAR-10 dataset.
@@ 7.1 Reduce Overfitting
As illustrated in Figure 4, each bar represents a specific training configuration: gray bars denote IID data distributions without augmentation, blue bars correspond to non-IID data distributions (Dirichlet $\alpha=0.1$), and orange bars indicate the use of data augmentation under IID settings.
Consistent with findings in prior inference attack literature, the success of topology inference attacks is closely tied to the degree of overfitting in the target model. To validate this, experiments were conducted on the CIFAR-10 dataset under four attack scenarios. By increasing the number of local training epochs from 3 to 10, the level of overfitting was elevated. Results show that, across all examined topologies, this change consistently improves the F1-Scores of attacks. This observation confirms that overfitting exacerbates the risk of topological information leakage in DFL systems.
To counteract overfitting, data augmentation techniques were employed, specifically random cropping with 4-pixel padding. The impact is evident in the substantial performance decline observed for the orange bars across all scenarios and topologies. This demonstrates that data augmentation not only enhances the generalization ability of local models but also effectively mitigates the success of topology inference attacks.
@@ 7.2 Data Heterogeneity
Furthermore, data distribution among clients plays a critical role in attack effectiveness. Non-IID data were simulated using a Dirichlet distribution with $\alpha=0.1$ to assess this. The results, represented by the blue bars, indicate a noticeable decline in attack performance across all scenarios, with the effect being most prominent in scenarios 3 and 4. This suggests that increased data heterogeneity introduces higher model divergence, which in turn hampers the inference model’s ability to exploit overfitted representations for edge prediction.
Overall, these findings imply that reducing overfitting through fewer local training epochs or data augmentation and increasing data heterogeneity are promising directions for mitigating topology inference risks in decentralized federated learning.
@@ 8 Conclusion and Future Work
This work introduces a novel topology inference attack against DFL, exposing critical vulnerabilities related to privacy and information leakage. By analyzing local models, attackers can accurately infer the overlay topology, one of DFL’s most sensitive assets, highlighting the system’s susceptibility to privacy breaches. The study explores various attack scenarios and develops tailored metrics, models, and algorithms, with experiments confirming the feasibility and effectiveness of these attacks. Furthermore, network size, density, model overfitting, and data heterogeneity significantly influence attack success. Mitigation strategies, such as data augmentation and increasing data heterogeneity, can enhance model generalization and reduce the risk of data leakage.
The present work focuses on attack scenarios 1 through 4 and does not yet address the most challenging attack scenario 5. Future research plans include exploring feasible strategies for Scenario 5 and extending the evaluations to a wider range of datasets and topologies. Besides, more attack metrics and strategies will be developed to improve attack effectiveness.